OpenStudy (anonymous):

how would I Identify the 20th term of the arithmetic sequence 3,7,11

OpenStudy (jdoe0001):

what's the common difference, "d", you think? that is, how much is being "added" to one term to get the next one?

OpenStudy (anonymous):

it's 4 @jdoe0001

OpenStudy (jdoe0001):

so d = 4 then \(\bf a_{\color{brown}{ n}}=a_1+({\color{brown}{ n}}-1){\color{blue}{ d}} \\ \quad \\ a_1=\textit{1st term}\qquad thus\quad a_{\color{brown}{ n}}=5+({\color{brown}{ n}}-1){\color{blue}{ (4)}} \\ \quad \\ 20^{th}\ term\implies a_{\color{brown}{ 20}}=5+({\color{brown}{ 20}}-1){\color{blue}{ (4)}}\)

OpenStudy (anonymous):

Lost me there

OpenStudy (anonymous):

@jdoe0001 I'm still confused

OpenStudy (jdoe0001):

http://www.youtube.com/watch?v=lj_X9JVSF8k <--- arithmetic sequences, which is what this one is

OpenStudy (anonymous):

@jdoe0001 its 79

OpenStudy (anonymous):

got it thanks i understand

OpenStudy (jdoe0001):

ohh yeah.. the 1st term was 3... yeap I had 5 as typo \(\bf a_{\color{brown}{ n}}=a_1+({\color{brown}{ n}}-1){\color{blue}{ d}} \\ \quad \\ a_1=\textit{1st term}\qquad thus\quad a_{\color{brown}{ n}}=3+({\color{brown}{ n}}-1){\color{blue}{ (4)}} \\ \quad \\ 20^{th}\ term\implies a_{\color{brown}{ 20}}=3+({\color{brown}{ 20}}-1){\color{blue}{ (4)}}\)

OpenStudy (anonymous):

but did i do it right @jdoe0001

OpenStudy (jdoe0001):

yeap \(\bf a_{\color{brown}{ n}}=a_1+({\color{brown}{ n}}-1){\color{blue}{ d}} \\ \quad \\ a_1=\textit{1st term}\qquad thus\quad a_{\color{brown}{ n}}=3+({\color{brown}{ n}}-1){\color{blue}{ (4)}} \\ \quad \\ 20^{th}\ term\implies a_{\color{brown}{ 20}}=3+({\color{brown}{ 20}}-1){\color{blue}{ (4)}}\implies a_{\color{brown}{ 20}}=3+19\cdot 4 \\ \quad \\ a_{\color{brown}{ 20}}=3+76\to 79\)